You have $7000 to invest for seven years. Which investment will provide the greater return, 5% compounded yearly or 4.75% compounded monthly?
Answer #1

P =$7000 - A=P( it ļjat t= 7 case : 1 Early {= 5%, compounded E-0.05 AzPC1+Eynt A = 7000 ( 1+ 0.05 ) ** A: 7000 (170.0537 4 =